# Conformal Prediction for Signal Temporal Logic Inference

### v1

@@ I Introduction
Contributions. Our contributions can be summarized as follows: (i) We design a new nonconformity score that operates directly on robustness values, even though ground‑truth robustness is unavailable, and develop smooth approximations that make this score suitable for training. (ii) We extend ConfTr based on (i) to handle STL inference models and provide a comparison with our proposed method. (iii) Leveraging inductive CP, we develop a smooth approximation of the CP procedure without explicit construction of prediction sets, allowing joint optimization of the inference model and its conformal wrapper; after training, non‑smoothed CP algorithm is applied to provide rigorous statistical guarantee on unseen data. (iv) We propose a novel loss function that combines the classification error and CP goal in a single term, unlike ConfTr, which relies on a separate regularization term. Our formulation requires no additional tuning and naturally avoids empty prediction sets. (v) In experiments, we show that our method improves both efficiency and coverage compared with state‑of‑the‑art conformal training baselines.
@@ II-A Signal Temporal Logic
STL Inference. Signal Temporal Logic Inference aims to learn an STL formula $\phi$ from a labeled dataset $\mathcal{S}:=\{(X^{(1)},Y^{(1)}),\cdots,(X^{(N)},Y^{(N)})\}$ with the label $Y^{(i)}\in\{1,-1\}$. Here $Y^{(i)}=1$ indicates that the signal $X^{(i)}$ exhibits a desired behavior, whereas $Y^{(i)}=-1$ indicates the opposite. During training, $\phi$ is treated as a classifier: for each $(X^{(i)},t)$, we predict $\tilde{Y}^{(i)}=1$ if $\rho^{\phi}(X^{(i)},t)>0$, otherwise we predict $\tilde{Y}^{(i)}=-1$. Equivalently, $(X^{(i)},t)\models\phi\iff\tilde{Y}^{(i)}=1$.
@@ IV-A STL margin
Definition 1 (Binary-class STL margin)
@@
Intuitively, eq. 8 defines the margin as the minimum distance from the origin to the robustness of correctly classified data points; this definition treats the distances for the positive and negative classes in a symmetric way. Specifically, as illustrated in Figure 1, the margin is given by $m_{\pi_{\theta},\mathcal{S}}=\min\{m_{p},m_{n}\}$. Thus, maximizing the margin helps to balance the decision boundary, preventing it from being biased toward one class, and thereby achieving a better separation between the two classes.
@@ IV-B Nonconformity Score
Proof:
Let $c>0$ represent a change of scale (i.e., a change of unit) from the robustness $\pi_{\theta}(X)$ to the robustness $\pi^{\prime}_{\theta}(X)=c\pi_{\theta}(X)$. From Definition 1, the corresponding margin is then $m^{\prime}=cm$. Multiplication by a positive constant preserves all inequalities in (10), so $\pi_{\theta}(X)\gtrless\pm m$ if and only if $\pi^{\prime}_{\theta}(X)\gtrless\pm m^{\prime}$; as a result, the assigned score is unchanged, i.e., $E_{\pi^{\prime}}(X,Y)=E_{\pi}(X,Y)$. ∎
@@ V-B ConfTr using differentiable thresholds
Calibration. A smooth quantile approximation, implemented with differentiable sorting techniques [25, 26] computes an approximated threshold $\tilde{\tau}$ being differentiable w.r.t. the model’s parameters $\theta$. This step is denoted as $\tilde{\tau}=\text{SMOOTHCAL}(\{\pi_{\theta}(X_{i}),Y_{i})\}_{i=1,...,n},\alpha)$.
Prediction. Given a test point $X$, the soft inclusion of a candidate label $k$ in the prediction set is
@@
We extend ConfTr to inference models by incorporating the differentiable nonconformity score from eq. 10, thereby extending conformal training to the temporal-logic setting.
@@ VI-A ConfTr with Inefficiency Regularization
where $\lambda\in\mathbb{R}$ adjusts the influence of the CP loss during training. The complete training procedure for ConfTr is outlined in Algorithm 1.
@@ VI-B TLICP with $p$-Value Optimization
A notable property of this loss is its independence from the confidence level $\alpha$. In ConfTr, $\alpha$ must be fixed during training to compute a threshold $\tau$, which ties optimization to that specific level and can halt progress once the predicted set reaches size one. Our $p$-value loss continues to encourage larger true-label $p$-values and smaller false-label $p$-values regardless of $\alpha$. Any desired $\alpha$ can then be selected post hoc with non-smooth CP, without retraining. Algorithm 2 summarizes the full procedure.
Remark 1
@@
Note that for hard examples where $p_{\theta}(X;Y)<\alpha$ and $p_{\theta}(X;-Y)>\alpha$, the corresponding terms in (17) are the same as (16); however, eq. 17 does not explicitly penalize easy examples.
@@ 2. Task 2
(a) VIMA simulated environment: the robot arm picks up a colored block from the table and places it into the designated basket region (red-white boundary) without violating the constraint (yellow bar).
(b) Solid lines indicate successful placements (endpoints inside the basket), while dashed lines indicate failures. The green box is the valid region (basket and constraint).
@@ VII-B Task 1: Placing a block into a basket
In the first part, we evaluate STL inference on trajectories for a pick-and-place task: place a block into the basket without violating state constraints. We compare TLICP with ConfTr and the Baseline. All methods use the nonconformity scores defined in Section IV-B. TLICP is independent of the confidence level $\alpha$, thus it is trained only once. ConfTr requires retraining for each different value of $\alpha$; in our case we train five models with $\alpha\in\{0.1,0.05,0.01,0.005,0.001\}$. After training, prediction sets are computed with non-smooth CP as the confidence level $\alpha$ varies from $0.001$ to $0.1$ (in steps of $0.001$). We report the inefficiency and misclassification rate (MCR); the results are shown in Figure 4 and Table Ia.
@@
Table Ia reports the MCRs, which are broadly similar across methods. Notably, at relatively larger $\alpha$ (0.1, 0.05), ConfTr’s MCRs are slightly higher than Baseline’s. This is likely because the weighted loss in (15): compared with the Baseline, which minimizes only classification error, ConfTr allocates capacity to reducing inefficiency (pushing it below 1) rather than MCR, which can hurt accuracy. For small $\alpha$, this tradeoff is less problematic because achieving size 1 is difficult thus shrinking the set generally also reduces MCR.
@@ VII-C Task 2: Placing two blocks into a basket in order
In the second part, we evaluate STL inference on a more complex task: placing block A and then block B into the basket without violating the constraint. We use the same setup as Task 1. Results are shown in Figure 4 and Table Ib.

### v3

# 1st Danyang Li    2nd Yixuan Wang    3rd Matthew Cleaveland    4th Mingyu Cai    5th Roberto Tron

@@ I Introduction
Contributions Our contributions can be summarized as follows: (i) We design a new nonconformity score that operates directly on robustness values, even though ground‑truth robustness is unavailable, and develop smooth approximations that make this score suitable for training. (ii) We extend ConfTr based on (i) to handle STL inference models and provide a comparison with our proposed method. (iii) Leveraging inductive CP, we develop a smooth approximation of the CP procedure without explicit construction of prediction sets, allowing joint optimization of the inference model and its conformal wrapper; after training, non‑smoothed CP algorithm is applied to provide rigorous statistical guarantee on unseen data. (iv) We propose a novel loss function that combines the classification error and CP goal in a single term, unlike ConfTr, which relies on a separate regularization term. Our formulation requires no additional tuning and naturally avoids empty prediction sets. (v) In experiments, we show that our method improves both efficiency and coverage compared with state‑of‑the‑art conformal training baselines.
@@ II-A Signal Temporal Logic
STL Inference Signal Temporal Logic Inference aims to learn an STL formula $\phi$ from a labeled dataset $\mathcal{S}:=\{(X^{(1)},Y^{(1)}),\cdots,(X^{(N)},Y^{(N)})\}$ with the label $Y^{(i)}\in\{1,-1\}$. Here $Y^{(i)}=1$ indicates that the signal $X^{(i)}$ exhibits a desired behavior, whereas $Y^{(i)}=-1$ indicates the opposite. During training, $\phi$ is treated as a classifier: for each $(X^{(i)},t)$, we predict $\tilde{Y}^{(i)}=1$ if $\rho^{\phi}(X^{(i)},t)>0$, otherwise we predict $\tilde{Y}^{(i)}=-1$. Equivalently, $(X^{(i)},t)\models\phi\iff\tilde{Y}^{(i)}=1$.
@@ IV-A STL margin
Definition 1 (Binary-class STL margin).
@@
Intuitively, Section 8 defines the margin as the minimum distance from the origin to the robustness of correctly classified data points; this definition treats the distances for the positive and negative classes in a symmetric way. Specifically, as illustrated in Figure 1, the margin is given by $m_{\pi_{\theta},\mathcal{S}}=\min\{m_{p},m_{n}\}$. Thus, maximizing the margin helps to balance the decision boundary, preventing it from being biased toward one class, and thereby achieving a better separation between the two classes.
@@ IV-B Nonconformity Score
Proof.
Let $c>0$ represent a change of scale (i.e., a change of unit) from the robustness $\pi_{\theta}(X)$ to the robustness $\pi^{\prime}_{\theta}(X)=c\pi_{\theta}(X)$. From Section 1, the corresponding margin is then $m^{\prime}=cm$. Multiplication by a positive constant preserves all inequalities in (9), so $\pi_{\theta}(X)\gtrless\pm m$ if and only if $\pi^{\prime}_{\theta}(X)\gtrless\pm m^{\prime}$; as a result, the assigned score is unchanged, i.e., $E_{\pi^{\prime}}(X,Y)=E_{\pi}(X,Y)$. ∎
@@ V-B ConfTr using differentiable thresholds
Calibration A smooth quantile approximation, implemented with differentiable sorting techniques [25, 26] computes an approximated threshold $\tilde{\tau}$ being differentiable w.r.t. the model’s parameters $\theta$. This step is denoted as $\tilde{\tau}=\text{SMOOTHCAL}(\{\pi_{\theta}(X_{i}),Y_{i})\}_{i=1,...,n},\alpha)$.
Prediction Given a test point $X$, the soft inclusion of a candidate label $k$ in the prediction set is
@@
We extend ConfTr to inference models by incorporating the differentiable nonconformity score from Section 10, thereby extending conformal training to the temporal-logic setting.
@@ VI-A ConfTr with Inefficiency Regularization
where $\lambda\in\mathbb{R}$ adjusts the influence of the CP loss during training. The complete training procedure for ConfTr is outlined in Section 1.
@@ VI-B TLICP with $p$-Value Optimization
A notable property of this loss is its independence from the confidence level $\alpha$. In ConfTr, $\alpha$ must be fixed during training to compute a threshold $\tau$, which ties optimization to that specific level and can halt progress once the predicted set reaches size one. Our $p$-value loss continues to encourage larger true-label $p$-values and smaller false-label $p$-values regardless of $\alpha$. Any desired $\alpha$ can then be selected post hoc with non-smooth CP, without retraining. Section 2 summarizes the full procedure.
Remark 1.
@@
Note that for hard examples where $p_{\theta}(X;Y)<\alpha$ and $p_{\theta}(X;-Y)>\alpha$, the corresponding terms in (17) are the same as (16); however, Section 17 does not explicitly penalize easy examples.
@@ VII-B Task 1: Placing a block into a basket
In the first part, we evaluate STL inference on trajectories for a pick-and-place task: place a block into the basket without violating state constraints. We compare TLICP with ConfTr and the Baseline. All methods use the nonconformity scores defined in Section IV-B. TLICP is independent of the confidence level $\alpha$, thus it is trained only once. ConfTr requires retraining for each different value of $\alpha$; in our case we train five models with $\alpha\in\{0.1,0.05,0.01,0.005,0.001\}$. After training, prediction sets are computed with non-smooth CP as the confidence level $\alpha$ varies from $0.001$ to $0.1$ (in steps of $0.001$). We report the inefficiency and misclassification rate (MCR); the results are shown in Figure 4 and Table LABEL:tab:task1_results.
@@
Table LABEL:tab:task1_results reports the MCRs, which are broadly similar across methods. Notably, at relatively larger $\alpha$ (0.1, 0.05), ConfTr’s MCRs are slightly higher than Baseline’s. This is likely because the weighted loss in (15): compared with the Baseline, which minimizes only classification error, ConfTr allocates capacity to reducing inefficiency (pushing it below 1) rather than MCR, which can hurt accuracy. For small $\alpha$, this tradeoff is less problematic because achieving size 1 is difficult thus shrinking the set generally also reduces MCR.
@@ VII-C Task 2: Placing two blocks into a basket in order
In the second part, we evaluate STL inference on a more complex task: placing block A and then block B into the basket without violating the constraint. We use the same setup as Task 1. Results are shown in Figure 4 and Table LABEL:tab:task2_results.